OpenStudy (nicole143):

@mathmale Hello! I had tried to ask someone else this question earlier but them didn't really explain it and I got more confused. I'm trying to do multiplication with square roots and variables. I will put up the equation, thank you if you can help(:

OpenStudy (mathmale):

Hello, N! I have time for one quick question. Go ahead.

OpenStudy (nicole143):

\[\sqrt{7x}\left( \sqrt{x} - 7\sqrt{7}\right) \]

OpenStudy (nicole143):

Oops *they :P

OpenStudy (mathmale):

What I'll do is to give you a quick review. But first, why the "oooops"?

OpenStudy (nicole143):

I had put "them" not "they" in the question ha

OpenStudy (mathmale):

\[\sqrt{7}\sqrt{7}=7,\sqrt{x}\sqrt{x}=x.\]

OpenStudy (mathmale):

\[\sqrt{7x)}=\sqrt{7}\sqrt{x}\]

OpenStudy (mathmale):

So, multiplying the first term within the parentheses by Sqrt(7x) is equivalent to multiplying it by Sqrt (7)*Sqrt(x). yOU OK WITH THAT?

OpenStudy (nicole143):

Yes, I believe so.

OpenStudy (mathmale):

\[\sqrt{7x}(\sqrt{x})=\sqrt{7}\sqrt{x}*\sqrt{x})=\sqrt{7}x\]

OpenStudy (mathmale):

How would you multiply the SECOND term within the parentheses by Sqrt(7x)? Please do this quickly on paper and then let me know your result.

OpenStudy (nicole143):

I think i would be; \[-7\sqrt{49x}\] ?

OpenStudy (mathmale):

If I were you I'd do this:\[\sqrt{7}\sqrt{x}*7*\sqrt{7}. Bet .you.can.finish.this.\]

OpenStudy (mathmale):

Try finishing mine; then compare our results.

OpenStudy (nicole143):

\[7\sqrt{49x}\] ??

OpenStudy (mathmale):

Please explain why that 49 goes under the radical sign.

OpenStudy (nicole143):

\[49\sqrt{x}\] Better?

OpenStudy (mathmale):

Much better. Actually, your answer is correct, except that it has not been reduced.

OpenStudy (mathmale):

So, your final result for the entire problem?

OpenStudy (nicole143):

What do I reduce it by?

OpenStudy (mathmale):

I don't think you need to reduce, not if you mean to reduce the entire answer.

OpenStudy (nicole143):

I don't know :/

OpenStudy (mathmale):

We started off with a problem that looked like a(b-c). First we multiplied the first term inside parentheses by a. Then we mult. the 2nd term inside par. by a. Now we need to combine the 2 products. The second product is the one we just got: -49Sqrt(x) If you'll go back to near the beginning you'll find the 1st product. (It's Sqrt(7) * x )

OpenStudy (mathmale):

Please write out the final solution now.

OpenStudy (nicole143):

\[\sqrt{7x} - 49\sqrt{x}\]

OpenStudy (mathmale):

Correct, EXCEPT that that first x should not be under the radical. Mind re-typing this for verification?

OpenStudy (nicole143):

\[\sqrt{7} x - 49\sqrt{x}\]

OpenStudy (mathmale):

Too nice for words! I need to quit. So glad to see you.

OpenStudy (nicole143):

Have a good night! Thank you! (:

OpenStudy (mathmale):

My pleasure! Ditto. :)